# Train Once for All: A Transitional Approach for Efficient Aspect Sentiment Triplet Extraction

### v1

@@ Abstract
Aspect-Opinion Pair Extraction (AOPE) and Aspect Sentiment Triplet Extraction (ASTE) have gained significant attention in natural language processing. However, most existing methods are a pipelined framework, which extracts aspects/opinions and identifies their relations separately, leading to a drawback of error propagation and high time complexity. Towards this problem, we propose a transition-based pipeline to mitigate token-level bias and capture position-aware aspect-opinion relations. With the use of a fused dataset and contrastive learning optimization, our model learns robust action patterns and can optimize separate subtasks jointly, often with linear-time complexity. The results show that our model achieves the best performance on both the ASTE and AOPE tasks, outperforming the state-of-the-art methods by at least 6.98% in the F1 measure. The code is available at https://github.com/Paparare/trans_aste.
Train Once for All: A Transitional Approach for Efficient Aspect Sentiment Triplet Extraction
Xinmeng Hou1  Lingyue Fu2  Chenhao Meng2  Hai Hu2thanks: Corresponding author 1Columbia University, New York, NY, USA 2Shanghai Jiao Tong University, Shanghai, China fh2450@tc.columbia.edu fulingyue@sjtu.edu.cn, chenhaomeng@sjtu.edu.cn, hu.hai@sjtu.edu.cn
@@ 1 Introduction
Aspect-Based Sentiment Analysis (ABSA) is a fine-grained sentiment analysis task that identifies specific aspects in text and analyzes the sentiments linked to them. It involves subtasks such as: Aspect Extraction (AE): Identifying mentioned aspects. Opinion Extraction (OE): Extracting opinion terms related to aspects. Aspect-Opinion Pair Extraction (AOPE): Extracting aspect-opinion pairs. Aspect-Sentiment Triplet Extraction (ASTE): Identifying aspect, opinion, and sentiment polarity triplets.
Figure 1: For the sentence ’Gourmet food is delicious. Good service, but not so welcoming,’ this example demonstrates the processing steps in both classic and transitional methods for extracting aspect-opinion pairs and tagging sentiment polarity. AE stands for Aspect Extraction, OE for Opinion Extraction, POS represents positive sentiment, and NEG represents negative sentiment.
Classic ABSA approaches often use a pipeline architecture, performing subtasks like aspect extraction (AE) and opinion extraction (OE) separately. For instance, in the sentence: “Gourmet food is delicious. Good service, but not so welcoming.,” AE identifies gourmet food and service as aspects, while OE extracts delicious, good, and not so welcoming as opinions. These outputs are combined to form aspect-opinion pairs, with sentiment tagging assigning polarities to create triplets. Past studies have generally achieved satisfactory performance on aspect and opinion extraction Jiang et al. (2023); Wang et al. (2023); Chakraborty (2024). Adding the according sentiment tags with pair, ASTE is the most integrated task for aspect-based sentiment analysis that has seen the development of diverse models leveraging various methodologies. Pipeline-based approaches, such as Peng-Two-stage Peng et al. (2020), decompose the task into multiple stages for modular refinement. Sequence-to-sequence frameworks like BARTABSA Yan et al. (2021) employ pretrained transformers to generate triplets flexibly. Sequence-tagging methods, including GTS Wu et al. (2020) and JET-BERT Xu et al. (2020), annotate tokens for precise identification of relationships. Machine Reading Comprehension (MRC)-based models, such as COM-MRC Zhai et al. (2022) and Triple-MRC Zou et al. (2024), reframe the task as query answering for efficient extraction. Graph-based approaches such as EMC-GCN Chen et al. (2022), BDTF Chen et al. (2022), and DGCNAP Li et al. (2023) use graph structures to capture semantic and syntactic interactions. Tagging schema-based models, exemplified by STAGE-3D Liang et al. (2023), use hierarchical schemas for multi-level extraction, while lightweight models like MiniConGTS Sun et al. (2024) focus on efficiency with reduced computational costs.
@@
Although previous studies have shown satisfactory results for individual entity extraction, performance on pair and triplet extraction remains suboptimal, with state-of-the-art (SOTA) scores in the 60s and 70s Sun et al. (2024). Two key challenges contribute to this gap: (1) Disconnected Aspect-Opinion Extraction: Opinions are often extracted independently from their corresponding aspects in previous studies Liang et al. (2023); Sun et al. (2024). While relationships can be added as an auxiliary factor to assist pair extraction Liu et al. (2022); Wang et al. (2023), this approach loses critical contextual information by treating aspects and opinions as separate entities. This limits the effectiveness of many token-based extraction methods. (2) Time Complexity with Longer Sequences: Methods using 2D matrix tagging Liang et al. (2023); Sun et al. (2024) to capture relationships between tokens face significant increases in time complexity as the length of the token sequence increases. This computational burden restricts their scalability, especially for longer texts in practical applications.
As an overview, the critical contributions of our study are as follows:
Efficient Transition Pipeline for AOPE:
We propose a transition pipeline for simultaneous aspect-opinion pair extraction, achieving linear time complexity $O(n)$.
Augmented Learning for Model Optimization:
Building on this transitional backbone, we introduce a contrastive-augmented optimization method to enhance model efficiency.
New SOTA Performance on AOPE and ASTE:
Our model achieves new state-of-the-art (SOTA) performance on both AOPE and ASTE tasks.
@@ 2 Related Work
2.1 Transition-based Methods
Transition-based approaches are widely used in dependency parsing, leveraging shift-reduce and bidirectional arc actions (left-arc, right-arc) for efficient $O(n)$ parsing Aho and Ullman (1973); Nivre (2003); Cer et al. (2010). These parsers maintain stack, buffer, and arc relations to track transitions and have been extended to tasks like token segmentation Zhang et al. (2016) and argument mining Bao et al. (2021). Beyond parsing, transition-based methods have been applied to structured sentiment analysis, generating graph structures for opinion extraction Fernández-González (2023) and incremental constituency parsing using graph neural networks Yang and Deng (2020). A unifying theory bridges transition-based parsing and sequence labeling for simplified learning Gómez-Rodríguez et al. (2020), while in AMR parsing, structure-aware fine-tuning achieves state-of-the-art results Zhou et al. (2021). Their flexibility and efficiency make them effective across diverse NLP tasks.
2.2 Contrastive-based Optimization
Contrastive learning has demonstrated its effectiveness in various domains, achieving state-of-the-art results in token-independent extraction tasks with models such as MINIconGTS Sun et al. (2024). Recent works have further explored its potential: contrastive learning for prompt-based few-shot language learners enhances generalization through augmented "views" Jian et al. (2022); contrastive learning as goal-conditioned reinforcement learning provides a novel loss function with theoretical guarantees for improved success in goal-directed tasks Eysenbach et al. (2023), and studies analyzing the role of margins reveal critical factors like positive sample emphasis for better generalization Rho et al. (2023). These advances showcase contrastive learning’s flexibility and robustness, which inspired us to integrate it into our transition-based AOPE and ASTE to refine representation learning and improve model performance.
3 Methodology
@@
Shift ($SF$) $(\sigma_{0},\,\beta_{0}\mid\beta_{1},E,C,R)\xrightarrow{SH}(\sigma_{0}\mid\sigma_{1},\beta_{1},E,C,R)$
Stop ($ST$) $(\sigma_{0},\O,A,O,R)\xrightarrow{ST}(\O,\O,A,O,R)$
@@
Right Constituent Removal ($R_{n}$) $(\sigma_{0}\mid\sigma_{1},\,\beta_{0},A,O,R)\xrightarrow{R_{n}}(\sigma_{1},\,\beta_{0},A,O,R)$
@@
Right-Relation Formation ($RR$) $(\sigma_{0}\mid\sigma_{1},\beta_{0},E,C,R)\xrightarrow{RR}(\sigma_{0}\mid\sigma_{1},\beta_{0},A\cup\sigma_{0},O\cup\sigma_{0},R\cup\sigma_{0}\xrightarrow{}\sigma_{1})$
Table 1: Symbolic Expressions for the Proposed Actions.
In this section, we introduce a novel paradigm that integrates the extraction of aspect-opinion pairs into a process reminiscent of constructing a parsing-directed graph. This model progressively constructs and annotates aspect-oriented relations using input sequences enriched with contextual features. Unlike the classic approach, which employs three sets to record action histories, we utilize five sets, including stack, buffer, aspect, and pair, to show how aspect and opinion are extracted sitimulately.
@@ 9 $ST$ [] [] [$\sigma_{1\&2}$] [$\sigma_{4}$] ($\sigma_{1\&2}\rightarrow\sigma_{4}$)
Table 1 provides a symbolic illustration of how the symbolic state is constructed and utilized. Take the sentence "Gourmet food is delicious" as an example. Table 2 demonstrates the process of moving tokens from the buffer to the stack, deciding whether they should be merged into a single entity or removed, and finally evaluating them for relation formation. It is important to note that the set of actions shown in the figure is not the only way to extract the "Gourmet food" and "delicious" aspect-opinion pair. The alternative solution for parsing the sentence is shown in Appendix 5.
@@ 3.3 Transition Implementation with Neural Model
This section outlines the transition-based parsing process using neural models. RoBERTa Liu et al. (2019) encodes textual phrases, while UniLSTM Hochreiter and Schmidhuber (1997) and BiLSTM Graves and Schmidhuber (2005) model dynamic transitions. The parser state is initialized and evolves through a sequence of actions with distributed embeddings and LSTMs representing the state at each step. At the end stage, a Multi-Layer Perceptron (MLP) is employed to classify the sentiment associated with each pair or triplet based on the evolved parser state.
@@ 3.4.1 Base Optimization
where $N$ is the number of samples, $M$ is the number of classes (either action or sentiment classes), $y_{i}^{j}$ is a binary indicator (0 or 1) indicating whether class $j$ is the correct class for sample $i$, and $p_{i}^{j}$ is the predicted probability for class $j$ for sample $i$. The total loss is the sum of the losses for both tasks:
$\mathcal{L}_{\text{base}}=\mathcal{L}_{\text{action}}+\mathcal{L}_{\text{sentiment}}.$ (6)
@@ 3.4.2 Contrastive-based Optimization
4 Experiments
4.1 Dataset
We used well-established benchmarks in Sentiment Analysis: SemEval 2014 (14lap, 14res) Pontiki et al. (2014), SemEval 2015 (15res) Pontiki et al. (2015), and SemEval 2016 (16res) Pontiki et al. (2016). These datasets, originally curated for the Semantic Evaluation (SemEval) competitions, are widely used for aspect-based sentiment analysis and extraction tasks Xu et al. (2021).
We employed a data processing framework to construct sentiment-enhanced dependency graphs for ASTE datasets. This framework utilizes SpaCy Honnibal et al. (2020) to extract syntactic dependencies and integrates sentiment scores from SenticNet Cambria et al. (2017) to generate weighted adjacency matrices. These enriched graph representations aim to expand the feature space and reduce biased learning. Along with tokenized sentences and their corresponding aspect-opinion-sentiment triplets, the processed data supports downstream tasks, including model training and evaluation.
Figure 3: F1 Score over epochs for the model trained on 14res, 14lap, 15res, and 16res datasets and tested on the 15res dataset. The results indicate that training on datasets other than 15res, such as 14res can yield an even better performance, which suggests the presence of richer action patterns in other datasets for the model to learn.
We evaluated the conformity of our transition operations in extracting aspects and opinions, with action evaluation results listed in Appendix A.3. Since our approach predicts actions rather than tokens, we assessed how well the model learns action patterns across individual datasets. For instance, as shown in Figure 3, the model did not perform best on the 15res test set when trained on the 15res training set. This suggests that other datasets, such as 14res, may contain richer action patterns for the model to learn. Consequently, we believe that training on a fused dataset provides a more diverse context, enhancing the model’s ability to learn these actions, which we then evaluated on separate test sets.
Figure 4: Growth on different weights for AOPE task on 14lap testing set.
Figure 5: Growth on different weights for ASTE task on 14lap testing set.

### v2

@@ Abstract
Aspect-Opinion Pair Extraction (AOPE) and Aspect Sentiment Triplet Extraction (ASTE) have drawn growing attention in NLP. However, most existing approaches extract aspects and opinions independently, optionally adding pairwise relations, often leading to error propagation and high time complexity. To address these challenges and being inspired by transition-based dependency parsing, we propose the first transition-based model for AOPE and ASTE that performs aspect and opinion extraction jointly, which also better captures position-aware aspect-opinion relations and mitigates entity-level bias. By integrating contrastive-augmented optimization, our model delivers more accurate action predictions and jointly optimizes separate subtasks in linear time. Extensive experiments on 4 commonly used ASTE/AOPE datasets show that, while performing worse when trained on a single dataset than some previous models, our model achieves the best performance on both ASTE and AOPE if trained on combined datasets, outperforming the strongest previous models in F1-measures (often by a large margin). We hypothesize that this is due to our model’s ability to learn transition actions from multiple datasets and domains. Our code is available at https://anonymous.4open.science/r/trans_aste-8FCF.
Train Once for All: A Transitional Approach for
Efficient Aspect Sentiment Triplet Extraction
@@ 1 Introduction
Aspect-Based Sentiment Analysis (ABSA) is a fine-grained sentiment analysis task that identifies specific aspects in text and analyzes the sentiments linked to them Hu and Liu (2004); Liu (2012); Wang et al. (2024). As shown in Figure 1, ABSA involves subtasks such as Aspect Extraction (AE) and Opinion Extraction (OE)—identifying mentioned aspects and their related opinions, or the combination—Aspect-Opinion Pair Extraction. Once the aspect and opinion have been extracted, a sentiment is usually computed, and this more complicated task is often referred to as Aspect-Sentiment Triplet Extraction (ASTE).
Figure 1: Demonstration of the processing steps in both classic and transitional methods for extracting aspect-opinion pairs and tagging sentiment polarity. Importantly, our proposed transitional method predicts transition actions, and performs pair extraction after the aspect–opinion relationship has been established, allowing the model to capture contextual relationships more effectively. AE stands for Aspect Extraction, OE for Opinion Extraction. POS and NEG represent positive and negative sentiments respectively.
For instance, given the sentence: “Gourmet food is delicious. Good service, but not so welcoming”, AE identifies gourmet food and service as aspects, while OE extracts delicious, good, and not so welcoming as opinions. These outputs are then combined to form aspect-opinion pairs, with a separate sentiment tagging system assigning polarities to create triplets Jiang et al. (2023); Wang et al. (2023); Chakraborty (2024).
@@
ASTE is the most integrated task for aspect-based sentiment analysis, for which diverse models leveraging various methodologies have been developed, including pipeline-based approach (Peng et al., 2020), sequence-to-sequence method (Yan et al., 2021), sequence-tagging method (Wu et al., 2020; Xu et al., 2020), to name just a few. Despite these efforts and growing interests, the accuracy of recent models remains suboptimal, with the best systems scoring 60% or 70% Sun et al. (2024). There are two key challenges that hinder performance: (1) Disconnected Aspect-Opinion Extraction: Opinions are often extracted independently from their corresponding aspects Liang et al. (2023); Sun et al. (2024). While positional relationships can be added as an auxiliary factor to assist pair extraction Liu et al. (2022); Wang et al. (2023), this approach loses critical contextual information by treating aspects and opinions as separate entities. This limits the effectiveness of many token-based extraction methods. (2) High time complexity with longer sequences: Methods using 2D matrix tagging Liang et al. (2023); Sun et al. (2024) to capture relationships between tokens face significant increases in time complexity as the length of the token sequence increases. This computational burden restricts their scalability, especially for longer texts in practical applications.
To address these two challenges, we present the first transition-based AOPE system named Trans-AOPE that (1) extracts the Aspect and the Opinion at the same time, and (2) has a time complexity of $O(n)$. We also introduce a contrastive-augmented optimization method to enhance model efficiency. We conduct experiments on 4 commonly used ABAS datasets, and compare our system with previous models. Our results show that Trans-model achieves state-of-the-art performance on all datasets we tested. We conduct comprehensive ablation studies to evaluate the contribution of optimization components and perform extensive training on various datasets to identify precisely where our model and baselines derive their learning. Our contributions are:
  • We propose the first transition-based model that extracts aspect-opinion pairs based on relational aspects, rather than using relational factors as supplementary references or confirmation, with linear time complexity.
  • We experiment with a contrastive-augmented optimization method and find that balanced weighting yields faster, more stable improvements, emerging as the optimal training configuration.
  • We explore various training strategies and show that our proposed method achieves optimal performance on four datasets when trained on combined training sets, with better cross-dataset generalization.
Method Approach Time Complexity
Peng-Two-stage (Peng et al., 2020) Two-Stage Pipeline: entity identification and relation formation $O(n+k^{2})$
BARTABSA (Yan et al., 2021) Generative-based Aspect-based Sentiment Analysis $O(m\cdot v)$
GTS (Wu et al., 2020) Grid Matrix-based Tagging $O(n^{2})$
JET-BERT Xu et al. (2020) Position-Aware Sequence Tagging $O(n)$
COM-MRC Zhai et al. (2022) Compositional Machine Reading Comprehension $O(r\cdot n^{2}\cdot h)$
Triple-MRC Zou et al. (2024) Multi-turn Machine Reading Comprehension $O(r\cdot n^{2}\cdot h)$
EMC-GCN Chen et al. (2022) Multi-channel Graph Convolutional Network $O(m\cdot n^{2}\cdot h)$
DGCNAP Li et al. (2023) Graph Convolutional Network w/ Affective Knowledge $O(m\cdot n^{2}\cdot h)$
MiniConGTS Sun et al. (2024) Lightweight Grid Matrix-based Tagging System $O(n^{2})$
Trans-model (Ours) Transition-based Action Prediction for Simulating Relation Formation and Pair Extraction $O(n)$
Table 1: An overview of previous methods and models (which will serve as baselines in this study), their approaches, and corresponding time complexities. Here, the hidden size for LSTM $d$ is simplified; $n$ is the sequence length; $m$ is the number of graph channels; $v$ is the vocabulary size; $k$ is the number of extracted terms; $r$ is the number of query rounds, and $h$ is the hidden size of the encoder.
@@ 2 Related Work
Previous methods on ASTE
Pipeline-based approaches, such as Peng-Two-stage Peng et al. (2020), decompose the task into multiple stages for modular refinement. Sequence-to-sequence frameworks like BARTABSA Yan et al. (2021) employ pretrained transformers to generate triplets flexibly. Sequence-tagging methods, including GTS Wu et al. (2020) and JET-BERT Xu et al. (2020), annotate tokens for precise identification of relationships. Machine Reading Comprehension (MRC)-based models, such as COM-MRC Zhai et al. (2022) and Triple-MRC Zou et al. (2024), reframe the task as query answering for efficient extraction. Graph-based approaches such as EMC-GCN Chen et al. (2022), BDTF Chen et al. (2022), and DGCNAP Li et al. (2023) use graph structures to capture semantic and syntactic interactions. Tagging schema-based models, exemplified by STAGE-3D Liang et al. (2023), use hierarchical schemas for multi-level extraction, while lightweight models like MiniConGTS Sun et al. (2024) focus on efficiency with reduced computational costs.
Table 1 summarizes these baseline methods, along with our proposed model, in terms of their core approaches and time complexities.
Transition-based Methods in NLP
Transition-based approaches are widely used in dependency parsing, leveraging shift-reduce and bidirectional arc actions (left-arc, right-arc) for efficient $O(n)$ parsing Aho and Ullman (1973); Nivre (2003); Cer et al. (2010). These parsers maintain stack, buffer, and arc relations to track transitions and then build up dependency relations between tokens.
Transition-based methods have also been applied to various NLP tasks, including token segmentation Zhang et al. (2016), argument mining Bao et al. (2021), constituency parsing Yang and Deng (2020), AMR parsing Zhou et al. (2021), and sequence labeling Gómez-Rodríguez et al. (2020), among others. Transition-based methods have been explored in emotion analysis Fan et al. (2020); Jian et al. (2024). In sentiment analysis, however, transition-based models have not been widely adopted. One exception is their use in generating graph structures for opinion extraction Fernández-González (2023), although this design relies on graph embeddings and thus results in a time complexity of $O(N^{2})$, with performance that lags behind more recent AOPE and ASTE approaches.
Contrastive-based Optimization
Contrastive learning has demonstrated its effectiveness in various domains, achieving state-of-the-art results in token-independent extraction tasks with models such as MiniconGTS Sun et al. (2024). Recent works have further explored its potential: contrastive learning for prompt-based few-shot language learners enhances generalization through augmented “view” Jian et al. (2022); contrastive learning as goal-conditioned reinforcement learning provides a novel loss function with theoretical guarantees for improved success in goal-directed tasks Eysenbach et al. (2023), and studies analyzing the role of margins reveal critical factors like positive sample emphasis for better generalization Rho et al. (2023). These advances showcase contrastive learning’s flexibility and robustness, which inspired us to integrate it into our transition-based AOPE and ASTE to refine representation learning and improve model performance.
@@
Shift ($SF$) $(\sigma_{0},\,\beta_{0}\mid\beta_{1},A,O,R)\xrightarrow{SH}(\sigma_{0}\mid\sigma_{1},\beta_{1},A,O,R)$
Stop ($ST$) $(\sigma_{0},,A,O,R)\xrightarrow{ST}(,,A,O,R)$
@@
Right Constituent Removal ($R_{n}$) $(\sigma_{0}\mid\sigma_{1},\,\beta_{0},A,O,R)\xrightarrow{R_{n}}(\sigma_{0},\,\beta_{0},A,O,R)$
@@
Right-Relation Formation ($RR$) $(\sigma_{0}\mid\sigma_{1},\beta_{0},A,O,R)\xrightarrow{RR}(\sigma_{0}\mid\sigma_{1},\beta_{0},A\cup\sigma_{0},O\cup\sigma_{1},R\cup\sigma_{0}\xrightarrow{}\sigma_{1})$
Table 2: Symbolic Expressions for the Proposed Actions. Here, $\sigma$ represents the stack, $\beta$ represents the buffer, $A$ denotes the aspect, $O$ denotes the opinion, and $R$ consist of an aspect and an opinion.
3 The Trans-AOPE/ASTE Model
In this section, we introduce a novel paradigm that treats the extraction of aspect-opinion pairs as a parsing-directed graph construction process. This paradigm comprises two major procedures: aspect-opinion pair extraction and tagging. We refer to the pair-extraction model as Trans-AOPE and the model that incorporates tagging as Trans-ASTE. Trans-AOPE progressively constructs and annotates aspect-oriented relations using input sequences enriched with contextual features. Unlike classic approaches, which employ three sets to record action histories, our paradigm utilizes five sets (including a stack, buffer, aspect set, and pair set) to demonstrate how aspects and opinions are extracted together.
@@ 9 $ST$ [] [] [$\sigma_{1\&2}$] [$\sigma_{4}$] ($\sigma_{1\&2}\rightarrow\sigma_{4}$)
Table 2 provides a symbolic illustration of how the symbolic state is constructed and utilized. Take the sentence "Gourmet food is delicious" as an example. Table 3 demonstrates the process of moving tokens from the buffer to the stack, deciding whether they should be merged into a single entity or removed, and finally evaluating them for relation formation. It is important to note that the set of actions shown in the figure is not the only way to extract the "Gourmet food" and "delicious" aspect-opinion pair. An alternative approach use the stack’s capacity of holding multiple tokens, moving "is" to the stack ($\beta_{3}\rightarrow\sigma_{3}$) before merging "Gourmet" and "food" ($[\sigma_{1},\sigma_{2},\sigma_{3}]\rightarrow[\sigma_{1\&2},\sigma_{3}]$).
@@ 3.3 Transition Implementation with Neural Model
This section introduces a transition-based parsing process. RoBERTa Liu et al. (2019) encodes the text, while UniLSTM Hochreiter and Schmidhuber (1997) and BiLSTM Graves and Schmidhuber (2005) capture transitions. The parser state evolves through a sequence of actions, with LSTMs processing each token once. This yields a time complexity of $O(n\cdot d^{2})$, typically simplified to $O(n)$ under fixed $d$. Finally, an MLP classifies the sentiment for each pair or triplet based on the final parser state.
@@ 3.4.1 Base Optimization
where $N$ is the number of samples, $M$ is the number of classes (either action or sentiment classes), $y_{i}^{j}$ is a binary indicator (0 or 1) indicating whether class $j$ is the correct class for sample $i$, and $p_{i}^{j}$ is the predicted probability for class $j$ for sample $i$. For AOPE task, total loss is action loss $\mathcal{L}_{\text{action}}$, and for ASTE task, is the sum of the losses for both tasks: $\mathcal{L}_{\text{base}}=\mathcal{L}_{\text{action}}+\mathcal{L}_{\text{sentiment}}.$
@@ 3.4.2 Contrastive-based Optimization
4 Experimental Setups
4.1 Datasets and preprocessing
We use well-established benchmarks in Sentiment Analysis: 14lap, 14res from SemEval-2014 Pontiki et al. (2014), 15res from SemEval-2015 Pontiki et al. (2015), and 16res from SemEval-2016 Pontiki et al. (2016) (see Table 4). The first is a collection of laptop reviews while the other three are curated from restaurant reviews. All are widely used for aspect-based sentiment analysis and extraction tasks Xu et al. (2021). Following previous literature, the four datasets are split in a 7:1.5:1.5 proportion for training, development, and test.
Datasets #S #POS #NEU #NEG #T
14res 2068 2869 286 754 3909
14lap 1453 1350 225 774 2349
15res 1075 1285 61 401 1747
16res 1393 1674 90 483 2247
Table 4: Statistics of four datasets. #S denotes the number data pieces, #POS, #NEU, #NEG the number of positive, neutral and negative sentiment labels, and #T the total number of triplets.
A data processing framework is used to construct sentiment-enhanced dependency graphs for ASTE datasets. This framework utilizes SpaCy Honnibal et al. (2020) to extract syntactic dependencies and integrates sentiment scores from SenticNet Cambria et al. (2017) to generate weighted adjacency matrices. These enriched graph representations aim to expand the feature space and reduce biased learning. Along with tokenized sentences and their corresponding aspect-opinion-sentiment triplets, the processed data supports downstream tasks, including model training and evaluation.
4.2 Training settings
We experiment with two settings for training data.
In-domain.
In this setting, we train with one of the four datasets and test on the test set of the same dataset, e.g., train on 14lap and test on 14lap.
Combined-train.
Since our method depends on the model learning the correct action to pair an aspect with an opinion from the training data, we hypothesize that it will have the advantage of being able to make use of training data from diverse domains to learn various actions. Thus we experiment with training on two or more training sets combined, to observe whether there is performance gain when more actions are learned. Specifically, we train on several training sets together, and evaluate on a single test set.